OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

Since you know y=2x, put 2x where y is in the other equation and solve for x. 2x=x+6

OpenStudy (anonymous):

Once you know x you can solve for y

OpenStudy (anonymous):

oh boy lol

OpenStudy (anonymous):

What did you get for x?

OpenStudy (anonymous):

4 i think

OpenStudy (anonymous):

3

OpenStudy (anonymous):

2x=x+6 x=6

OpenStudy (anonymous):

Now since you know x=6 put that into one of the equations and solve for y

OpenStudy (anonymous):

i hate math

OpenStudy (anonymous):

y=2x put 6 where x is y=2(6) y=12 so your answer is (6,12)

OpenStudy (anonymous):

dang lol

OpenStudy (anonymous):

man thank u so much

OpenStudy (anonymous):

you're welcome